In a stationery store, pencils have one price and pens have another price. Two pencils and three pens cost 78 cents. But three pencils and two pens cost 72 cents. How much does one pencil cost?

Answer by mananth(16946)

let cost of pencil be x
let cost of pen be y
2x+3y=78-------------1
3x+2y=72-------------2
multiply eq1 by 3 & eq 2 by 2
6x+9y=234
6x+4y=144
solve for y by subtracting the equations
5y=90
y=90/5=18
substitute the value of y in eq1
2x+3*18=78
2x=78-54
2x=24
x= 12
pencil costs 12 cents and pen costs 18 cents